The following gives the geometric sequence 2,M,N,250 find the values of M and N
Answer by Edwin McCravy(20059)

 2,       M,      N ,     250  
 ∥        ∥       ∥        ∥  
 2,      2r,     2r²,     2r³
 ∥        ∥       ∥        ∥ 
2=2,    M=2r,   N=2r²,  2r³=250
          ∥       ∥      r³=125      
          ∥       ∥       r=5
          ∥       ∥        
       M=2r,    N=2r²
       M=2(5)   N=2(5)²
       M=10     N=2(25)
                N=50 

Sequence: 2,M,N,250 = 2,10,50,250 
 
Edwin